-|b+4|=10 what is the solution?
Answer by jim_thompson5910(35256)

-|b+4|=10 is the same as |b+4| = -10 after multiplying both sides by -1

There are NO solutions to either equation because |b+4| = -10 has no solution. The absolute value of any number is never negative.